OpenStudy (anonymous):

Carbon dating. How old is an ivory tusk that has lost 40% of its carbon-14?

OpenStudy (wolf1728):

You need the half-life of C-14 which is 5730 years.

OpenStudy (wolf1728):

The formula is elapsed time = 5,730* log [(begng amt)/(ending amt)]/ log 2 time = 5,730 * log (100%/40%) log 2 time = 5,730 * log 2.5 / log 2 time = 5,730 * 0.3979400087 / 0.3010299957 time = 7,575 years and a calculator to check your answer http://www.1728.org/halflife.htm

OpenStudy (anonymous):

Oh I see. Thanks that helped out alot

OpenStudy (wolf1728):

Hmmmmmmmm LOST 40% so that means 60% remains. have to recalculate 5,730 * log(100%/60%) / log2

OpenStudy (wolf1728):

5,730 * log(1.6666666666) / log 2 5,730 * 0.2218487496 / 0.3010299957 =4,222.8 years or rounded 4,223 years